Mixture Problem:
Question: Determine how many liters of 6% acid solution and how many liters of 12% acid solution should be mixed by a chemist to obtain 6 liter of 10% acid solution.

Ans. 3.6 ---> but i dont know the solution..

Answer by ankor@dixie-net.com(22740)

Determine how many liters of 6% acid solution and how many liters of 12% acid solution should be mixed by a chemist to obtain 6 liters of 10% acid solution.
:
Let x = amt of 12% solution
the resulting amt is to be 6 liters, therefore:
(6-x) = amt of 6% solution
:
A typical mixture equation, using the the decimal equiv of per cent
.12x + .06(6-x) = .10(6)
.12x + .36 - .06x = .6
.12x - .06x = .6 - .36
.06x = .24
x = .24/.06
x = 4 liters of the 12% solution
and
6 - 4 = 2 liters of 6% solution
:
:
Confirm this solution
.12(4) + .06(2) = .10(6)
.48 + .12 = .6;
I don't know how they got a solution of 3.6 from this; anyway I wish you and your family a Happy Thanksgiving!!!